Solve the inequalities. Give your answer in interval notation, and indicate the answer geometrically on the real-number line. a. t + 6 ≤ 2 + 3t

b. 3(2 – 3x) > 4(1 – 4x) 


1
Expert's answer
2022-05-22T23:51:38-0400

a.


"t + 6\le 2 + 3t"

"3t-t\ge6-2"

"2t\ge4"

"t\ge 2"

"t\in [2, \infin)"




b.


"3(2 \u2013 3x) > 4(1 \u2013 4x)"

"6-9x>4-16x"

"6-9x+16x-6>4-16x+16x-6"

"7x>-2"

"x>-\dfrac{2}{7}"

"x\in (-\dfrac{2}{7}, \infin)"